OpenStudy (anonymous):

Use Cramer's rule to solve the system of equations. {x-2y=-9 3x+y=1

OpenStudy (amistre64):

\[D,Dx,Dy\] \[7,28,7\] x = -Dx/D y = Dy/D

OpenStudy (amistre64):

got me Dx and Dy bacwards :)

OpenStudy (amistre64):

7,7,28

OpenStudy (amistre64):

\[D,Dy,Dx\] \[ \begin{pmatrix}1&-2\\3&1\end{pmatrix}_{1}^{(-6)}, \begin{pmatrix}1&-9\\3&1\end{pmatrix}_{1}^{(-27)}, \begin{pmatrix}-9&-2\\1&1\end{pmatrix}_{-9}^{(-2)} \] \[7,28,-7\] should be it ...

OpenStudy (amistre64):

y= 28/7 = 4 x = -7/7 = -1 x - 2y =-9 -1 4 ---------- -1 - 9 = -9 3x +y=1 -1 4 -------- -3+4 = 1

OpenStudy (amistre64):

works for me :)